OpenStudy (anonymous):

Kathryn currently has an account balance of $2,248.95. She opened the account 7 years ago with a deposit of $1,762.14. If the interest compounds quarterly, what is the interest rate on the account?

OpenStudy (anonymous):

help

OpenStudy (amistre64):

work the math .. A = P(1+r/n)^(nt) solve for r

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\to \$2,248.95\\ P=\textit{original amount deposited}\to \$1,762.14\\ {\color{blue}{ r}}=rate\to (100\cdot r)\%\\ n=\textit{compounding cycle, Quaterly, means 4 times per year}\to 4\\ t=years\to 7 \\ \quad \\ 2,248.95=1,762.14\left(1+\frac{{\color{blue}{ r}}}{4}\right)^{4\cdot 7}\) as amistre64 suggested, solve for "r"

OpenStudy (anonymous):

am confuse

OpenStudy (jdoe0001):

hmm ? that's the compound interest formula...above.... there's an assumption you've covered it already

OpenStudy (anonymous):

i've got 3.5 percent

OpenStudy (jdoe0001):

hmm .. how .... hmmm how did you get 3.5%?

OpenStudy (jdoe0001):

what did you end up with "r" as anyhow? r= ?

OpenStudy (anonymous):

i got it thk you

OpenStudy (jdoe0001):

yw